A 22-foot rope is cut into 2 pieces. Seven less than four times the length of the longer rope piece x is five times the length of the shorter piece, which is (22 − x) ft. Find the length of the longer rope piece x (in ft).
Answer by Boreal(15235)

seven less than 4 times something is 4x-7, if x is the "something".
4x-7=5(22-x)
4x-7=110-5x
9x=117
x=13 ft.
That is the longer part of the rope
the shorter is 22-x, or 9 feet.
4x-7=45.
5(22-x)=5(9)=45.
This checks.
The shorter piece is 9 ft, the longer 13 ft.